OpenStudy (anonymous):

-3x^4+27x^2+1200=0 find all zeros of the equation!

OpenStudy (anonymous):

@Hero PLEASE HELP D: ive been on this for 3 hours

OpenStudy (anonymous):

Please show WORK. NO GRAPHING. I NEED WORK. I have some, but i cant get the correct answer.

OpenStudy (a1234):

First, pull out the GCF. -3x^4+27x^2+1200=0 -3(x^4 - 9x^2 - 400) = 0 Find two factors that have a product of 400 and a sum of 9. -3(x^4 - 25x^2 + 16x^2 - 400) = 0 -3(x^2(x^2 - 25) + 16(x^2 - 25)) = 0 -3(x^2 + 16)(x^2 - 25) = 0 -3(x^2 + 16)(x + 5)(x - 5) = 0 (x^2 + 16) ---> x^2 = -16 ---> this doesn't count because sqrt -16 cannot be done. (x + 5) ---> x = -5 (x - 5) ---> x = 5 The zeros are x = 5 and x = -5.

OpenStudy (anonymous):

also 4i and -4i?

OpenStudy (a1234):

Personally, I've never been asked for those solutions when I've done these types of problems.

OpenStudy (anonymous):

oh ok well u did good! Can i ask u a different question?????

OpenStudy (a1234):

If it's something similar, yes.

OpenStudy (anonymous):

What is the equation of y=x^3 with the given transformations? vertical stretch by a factor of 3,horizontal shift 4 units to the right, vertical shift 3 units down.

OpenStudy (a1234):

It doesn't seem like I know how to do that... What grade is this? Just curious.

OpenStudy (anonymous):

10 algebra 2

OpenStudy (a1234):

Well, I'm doing Algebra 1 :).

OpenStudy (anonymous):

oh wow